# Proximity and Visuotactile Point Cloud Fusion for Contact Patches in Extreme Deformation

### v1

@@ V Experimental Methods
TABLE I: Parameters for ground truth method.
Object Membrane Temperature (Celsius) Contact Time (s)
@@ V-A Ground Truth
It has been challenging to provide an experimental ground truth method for contact patches because the contact patch is always occluded from external viewpoints and surface mounted sensors interfere with the object interaction. Thus, most works use simulation or model-based methods [41, 35, 42] to evaluate the contact patch estimation. However, our work targets scenarios when simulation or model-based methods fail, so we must use an experimental method for ground truth.
Our ground truth method leverages thermal conduction at the contact interface. Similar methods have been presented to annotate human grasps on objects by capturing heat transfer from human hands with either thermal cameras [43] or thermochromic paint [44] and RGB-D cameras.
The object dataset is painted with three coats of thermochromic paint (Elmers), which changes color from purple (cold) to pink (hot) with temperature. The objects are placed in a fridge and cooled to ${}^{\circ}$ until the purple color is opaque and visually consistent, usually 15 minutes. Prior to running the experiment, the membrane is heated to temperatures ranging from ${}^{\circ}$${}^{\circ}$ with a heat gun. The precise temperature of the membrane and contact time is tuned per object to prevent noise from thermal radiation; the chosen temperature prevents color change when the object is within 1 mm of the membrane, but enables color change when contact occurs. The temperatures and contact time vary per object because each object has a different specific heat capacity, which governs the temperature at which the thermochromic paint changes color (Table I). We monitor the spatial uniformity of the membrane and object temperature with a FLIR thermal camera.
When the cooled object makes contact with the heated membrane, the thermochromic paint changes color where contact occurs. The object is then scanned with a high-resolution 3D scanner (Space Spider, Artec), with an accuracy of 0.05mm. The 3D scan of the object is then thresholded by color to determine the contact patch geometry.
We make the following two assumptions for this ground truth method:
  1. 1.
    Because all of the objects are relatively more rigid than the soft membrane, we assume that the object does not deform when contact occurs, so the scan of the object geometry is equivalent to the contact patch geometry.
  2. 2.
    This method captures the maximum contact patch, which we assume to occur when the membrane is at its highest deformation during its interaction with the object.
@@ V-B Membrane Strain Measurement
We use the following method to correlate displacement of the object into the membrane to 2D strain in the membrane (Figure 5). High strain in the membrane is difficult to simulate accurately due to its hyperelastic material properties and manufacturing imperfections, so we take an experimental approach to strain measurement.
First, we use a laser-cut stencil to pattern an IR and visibly opaque membrane with a uniform grid of dots with known spacing (10mm). This membrane is fabricated to have the same thickness and physical dimensions as the membrane described in Section IIIA. We use the Realsense L515 to capture an RGB-D point cloud of the patterned membrane as the robot arm applies a vertical displacement with an object from our object dataset. We measure the distances between the centroids of each dot on the patterned membrane from the point cloud, which is used to calculate the maximum in-plane strain along the length and width of the membrane. The calculated maximum strain is then used to find the vertical displacement required to impart the desired strain states for our experiments: low - 10%, medium - 60%, high - 100%.
@@ V-C Baselines
Figure 6: This membrane is opaque to IR and visible light. It is used for the tactile-only and membrane mechanics model baselines.
For the tactile-only and membrane mechanics model baselines, we use an unpatterned IR-opaque silicone membrane with the same physical dimensions as the selectively transmissive membrane (Figure 6). We use the Realsense L515 ToF depth camera to collect data for both of these baselines, since using a ToF depth camera with this sensor system can replicate a closely related design [8] and fulfill the input requirement of ToF depth data for the membrane mechanics model. For the proximity-only baseline, we use the selectively transmissive membrane and the depth data from the Realsense L515 ToF camera. The computation times for each algorithm are shown in Table 2.
@@ V-C2 Proximity-Only
The proximity-only approach exclusively uses the proximity point cloud. This baseline uses a reference point cloud of the inflated membrane with no object-induced deformation. The contact patch is estimated as any points that are closer to the camera than the reference point cloud.
@@ V-C3 Membrane Mechanics Model
TABLE II: Computation time for one frame of a contact patch estimation, using an Intel Core i7-8650U CPU.
@@
tactile-only (baseline) 0.03
proximity-only (baseline) 0.03
@@
mechanics model (baseline) 0.3
@@ VI Evaluation
To evaluate our proposed algorithm, we measure the accuracy of the geometry of the estimated contact patch when the membrane is in low (approx. 10%), medium (approx. 60%), and high (approx. 100%+) strain states. We use a dataset of three objects: a stuffed octopus, a Rubiks cube, and a cup. These objects were chosen to represent a diverse set of geometries: edges (Rubiks cube), curves (octopus), and non-convex contacts (cup). Each object is grasped by a Robotiq gripper mounted to a UR10 robot arm and pushed into the sensor. The vertical displacement of each trial is determined by what is needed to induce the desired strain in the membrane.
We compare our results to ground truth and three baselines: tactile-only, proximity-only, and the mechanics model-based method. The metric for geometric accuracy of the contact patch point clouds is root mean squared error (RMSE) of the estimated contact patch point cloud and the ground truth point cloud after manual alignment. For the membrane mechanics model-based method, we crop the estimated contact patch to exclude the outer edges of the sensor interface since the model cannot consistently predict the “bulging” behavior of the membrane, leading to spurious estimated contacts detected there. The resulting contact patch estimations are shown in Figure 7 and the RMSE analysis per strain state is shown in Figure 8.
@@
Figure 8: RMSE for the contact patch point clouds estimated by the fusion (ours), proximity only, tactile only, and membrane mechanics model algorithms per object. Lower RMSE is better.
Figure 9: Average RMSE for the contact patch point clouds estimated by the fusion (ours), proximity only, tactile only, and membrane mechanics model algorithms per strain state. Lower RMSE is better.
VI-A Low Strain Experiments
We induce low strain in the membrane by commanding the UR10 robot arm to press the object 4 mm into the membrane. The initial maximum height of the membrane is 314 mm from the internal cameras at the base of the sensor.
For the octopus, the contact patch estimation from the fusion algorithm has the lowest RMSE (Figure 8). Visually, the geometry of the contact patch estimated by the fusion algorithm also best matches the geometry of the ground truth patch (Figure 7). The proximity-only algorithm overestimates the contact patch. The overestimation of the contact patch is because the proximity-only algorithm does not consider the behavior of the membrane after contact, so it is unable to estimate where the object is actually in contact with the membrane. However, the geometry of the contact patch is still quite accurate because the proximity modality can measure the object geometry accurately. The tactile-only algorithm underestimates the contact patch, likely because the threshold is not specifically tuned to this strain range. At a low strain range, the threshold should be increased to capture more of the deformed points, since more of the membrane deformation can be attributed to the contact itself. However, increasing this threshold number may not generalize well to higher strain states, as the membrane starts to deform beyond the object’s contact patch. The membrane mechanics model estimates a very small patch, with the geometry of the patch largely dictated by the triangular elements of the mesh.
The fusion algorithm also performed the best compared to the baseline methods for the cup with handle, with separate multiple contacts. Some noise from IR reflections on the surface of the membrane was not completely eliminated by the RGB masking in the proximity modality, resulting in an extraneous contact patch detected. Using the tactile sensing modality helped make the extraneous contact patch smaller, but did not completely remove it because of the imperfect height alignment of the internal depth cameras and the tolerance used to compute the intersection of the point clouds. The extraneous contact patch is much more prominent in the proximity modality, which produced the third-best estimation of the contact. The estimated contact from tactile sensing is the second-best estimation. Compared to ground truth, it is difficult to visually distinguish which parts of the estimated patch are actually in contact with the cup, due to the noise from the membrane deformation propagating beyond the contact patch. The membrane mechanics model failed to produce an estimate of the contact patch, because the deformation is within the same magnitude as the simulated depth camera noise and thus indistinguishable.
With the cube, the tactile-only, proximity-only, and fusion algorithms all performed similarly and very well, with RMSEs less than 2 mm and within 16% of each other. The flat, planar face of the cube produces a large and dense contact patch with the membrane and the edges of the cube are captured very well with the proximity-only and fusion algorithms. In comparison, the tactile-only contact patch has significantly more noise around the border of the cube. The membrane mechanics model failed to estimate the contact patch, for the same reason as the cup.
Figure 10: We demonstrate the proposed fusion algorithm for contact patch estimation in two applications with complex membrane mechanics: a varied stiffness membrane (top) and membrane wrinkles (bottom).
VI-B Medium Strain Experiments
We program the robot arm to press the object 24 mm into the membrane for the membrane’s medium strain state. Similar to the low strain experiments, the initial maximum height of the membrane is 314 mm.
For the octopus, the fusion algorithm had the best performance compared to the baselines. Similar to the estimations in the low strain experiment, the proximity-only baseline overestimated the contact patch while the tactile-only baseline underestimated the contact patch. Compared to the tactile-only baseline, the proximity-only baseline and fusion algorithm better captured the geometry of the outline of the ground truth contact patch. The membrane mechanics model greatly underestimated the contact patch, with the overall geometry of the patch dictated by the coarse mesh of the membrane - the triangular edges are clearly seen in the geometry of the contact patch estimation.
For the cup, the fusion algorithm clearly best estimates the contact patch. Here, the advantages of the fusion approach is most clearly demonstrated: the tactile-only baseline is not able to distinguish two different contacts from the edge of the cup and the handle, while the proximity-only baseline estimates significant extraneous contact along the side of the cup near the handle. However, the fusion algorithm clearly distinguishes the two contacts while avoiding the same overestimation as proximity. Similar to the octopus, the membrane model estimation is again dominated by the geometry of the coarse mesh of the membrane, and also is unable to distinguish two different contact points of the cup base and handle. Like in the small strain case, it is also possible that the deformation from the cup handle is within the same magnitude as the simulated depth camera noise and the mechanics model failed to produce an estimate of the cup handle contact.
With the cube, the tactile-only baseline and fusion algorithms both produced similarly accurate estimates of the contact patch. The proximity-only baseline greatly overestimates the contact patch, estimating substantial contact along the sides of the cube. The mechanics model significantly underestimates the size of the contact patch, but is able to roughly capture the square geometry of the cube.
VI-C High Strain Experiments
The membrane is inflated to have an initial maximum height of 380 mm. The UR10 robot arm presses the object 120 mm into the membrane to produce the high strain state of the membrane. The object was pressed off-center into the membrane to allow the uncontacted region to expand outwards and produce high strains in the membrane.
The proximity-only baseline provided the most accurate contact patch for the octopus, despite slightly overestimating the size of the patch. The fusion and tactile-only baseline performed very similarly and both underestimated the patch. The mechanics model did produce a patch estimate but did not capture the circular geometry of the object well, estimating distinct corners along the edges of the patch.
For the cup, the fusion algorithm again excels capturing the two distinctive contact patches from the base of the cup and the handle. The proximity-only baseline greatly overestimates the contact patch, with more noticeable error along the sides of the cup. The mechanics model significantly underestimates the contact patch and is not able to distinguish the contact from the handle of the cup. The tactile-only baseline is not able to distinguish separate contacts between the handle and the base of the cup. In this case, because the cup is pressed off-center into the membrane, the tactile-only threshold approach shows significant bias in overestimating contact in the area of the membrane that undergoes high strain as it bulges towards the center. Because the threshold approach is only based on identifying total deformation, it will always be biased towards estimating false contacts along areas of high strain, even if contact did not occur.
The proximity-only and tactile-only baseline contact estimations for the cube also show the same limitations of thresholding in an off-center contact case, and overestimate contact in areas where the membrane is experiencing high strain. We note that the proximity-only estimation of the cube likely includes noise from IR light reflecting off the membrane on the right, since it is impossible to see two sides of a cube simultaneously. The fusion algorithm produces the most accurate estimation, although it underestimates the amount of contact along the side of the cube that is closest to the center of the sensor. The mechanics model does produce a contact patch estimate in approximately the correct location, but it does not resemble the geometry of the contact patch.
VI-D Discussion
@@
Although our proposed fusion algorithm is simple, it generalizes well across our object dataset and strain range, outperforming all of the baselines on average (Figure 7). Fusing both the tactile and proximity modalities is key to overcoming the shortcomings of each modality: the proximity modality tends to overestimate contact while tactile thresholding tends to underestimate contact. There are only two specific cases where a baseline has a lower RMSE than the fused contact patch: the proximity-only estimations in the high-strain octopus contact patch and the low-strain cube patch. With both of these cases, the proximity-only baseline overestimation of the contact patch better captured the ground-truth contact patch geometry than the underestimation from the fusion algorithm. In all other cases, the fusion algorithm consistently produced the most accurate contact patch with the lowest RMSE error.
@@ VII Demonstrations
We demonstrate several applications of the contact patch fusion algorithm to highlight the usefulness of fast and accurate contact patch estimation during complex membrane deformations: varied stiffness membranes, torque and shear-induced membrane wrinkling, closed loop control, and pose estimation.
@@ VII-A1 Varied Stiffness Membranes
Spatially varied stiffness patterns can be used to program complex 3D shapes [45, 46] that could be tailored towards manipulation tasks, such as directing the inflation of the membrane to a flatter shape for larger contact areas with less force, rather than a uniform and curved expansion of the membrane. Additionally, actively controlled variable stiffness materials [47, 48] are a rich area of soft robotics research and could potentially be used to change the stiffnesses and shapes of robot end effectors to suit different tasks in real time. Because the selectively transmissive membrane exclusively relies on optical properties to enable the tactile and proximity modalities, the stiffness of the membrane can be varied with no impact to the sensing functionality. This sensing approach, used in conjunction with the mechanics-independent proximity and tactile fusion contact patch algorithm, enables contact patch estimation for arbitrary patterns of varied stiffness in the sensing membrane.
In this demonstration, we achieve varied stiffness by using two different base silicone elastomers with different Shore hardnesses (Figure 8). The stiff and soft silicones are cured together to form one cohesive varied stiffness membrane. The two stiff regions are made from a stiffer silicone, Dragonskin 10 Fast (Smooth-On), which has a Shore hardness of 10A. The two soft regions are made from Ecoflex 00-35 Fast (Smooth-On), which has a Shore hardness of 00-35. Dragonskin 10 is approximately 8x harder than Ecoflex 00-35. The same fabrication method presented in Section IIIA is used to produce the varied stiffness membrane, albeit with the above specified base elastomers.
Here, the advantage of our approach is the ease-of-use across many different stiffness patterns and potentially actively controlled variable stiffness membranes. The proximity-only and tactile-only baseline methods would require tuning for each membrane pattern due to the variance of inflated shapes. The membrane mechanics model would require generating a new mesh that matches both the stiffness pattern and mechanical properties of the different membrane patterns. A machine learning approach would require more data for each new membrane pattern.
We run the medium strain cube experiment with the varied stiffness membrane. The estimated contact patch is shown in Figure 8. Compared to ground truth, the RMSE of the estimated contact patch is 1.09 mm.
@@ VII-A2 Torque and Shear-Induced Wrinkling
A highly deformable contact interface that tolerates a large range of shear and torque before slipping is key to maintaining stiction with a target object. This feature can be critical in mitigating errors in robot manipulators by reducing the requirement for precise perception and control during a task, or simplifying planning by minimizing regrasping or repositioning the object. However, these torques and shear forces can create wrinkles in the deformable membrane surface, which are very difficult to model even with computationally expensive FEM approaches. In this demonstration, we highlight the advantage of our mechanics-independent, fusion contact patch algorithm to estimate the contact patch of a cup while the membrane is wrinkled. The robot arm presses the cup into the surface of the sensor and rotates it ${}^{\circ}$, creating wrinkles behind the cup handle and along the sides of the cup. The estimated contact patch is shown in Figure 8. In comparison with ground truth, the RMSE of the estimated contact patch is 3.49 mm. The geometry of the contact patch qualitatively matches well with ground truth, accurately capturing the side of the cup handle that is in contact with the membrane despite the wrinkles.
@@ VII-B Closed-Loop Control for Whole-Body Manipulation
We demonstrate the real-time speed and integration of the proximity and tactile fusion algorithm with a closed-loop controller and physical robot system for a whole body manipulation task: balancing a stack of books (Figure 9). The robot system, Punyo-1 [49], consists of the visuotactile and proximity sensor placed on the robot’s torso, a Kinova Jaco robot arm on each side of the sensor with soft padding, and a Puget computer. An empty cardboard tray is held by Punyo-1 and pressed into the sensorized torso. As each book is placed on the tray held by the robot, the tray’s center of mass changes and the robot adjusts the roll angle of the tray to balance the books. The angle of the tray of books is sensed with the proximity and visuotactile fusion contact patch algorithm.
The sensor runs at 30 FPS and the contact patch is calculated for each frame. To calculate the angle of the tray, we use principal component analysis (PCA) on the contact patch. We found that PCA of the contact patch gives less noisy estimations of the tray position compared to the proximity-only modality, which has added noise from sensing the books. The estimated angle of the tray is then given to the whole-body PD controller at a frequency of 29 Hz. We use two arm motion primitives to adjust the angle of the tray.

### v2

@@ V Experimental Methods
Object Membrane Temperature (C) Contact Time (s)
@@ V-A Ground Truth
Our ground truth method leverages thermal conduction at the contact interface, similar to [25, 26] (Fig. 2).
We paint the object dataset with three coats of thermochromic paint (Elmers), which changes color from purple (cold) to pink (hot) with temperature. The objects are cooled to 3.5C until the purple color is opaque. The membrane is then heated to 70C-100C with a heat gun. The precise temperature of the membrane and contact time is tuned per object to prevent noise from thermal radiation; the chosen temperature prevents color change when the object is within 1 mm of the membrane, but enables color change when contact occurs. The temperatures and contact time vary per object because each object has a different specific heat capacity (Table I). We monitor the spatial uniformity of the membrane and object temperature with a FLIR thermal camera.
The thermochromic paint changes color where contact occurs. We scan the object with a high-resolution 3D scanner (Space Spider, Artec) and threshold the scan by color to produce the ground-truth contact patch. This method captures the maximum contact patch, which we assume to occur when the membrane is at its highest deformation during its interaction with the object.
@@ V-B Membrane Strain Measurement
We use the following method to correlate displacement of the object into the membrane to 2D strain in the membrane (Figure 3). First, we use a laser-cut stencil to pattern an IR and visibly opaque membrane with a uniform grid of dots with known spacing (10mm). This membrane is fabricated to have the same thickness and physical dimensions as our sensing membrane. We use the Realsense L515 to capture an RGB-D point cloud of the patterned membrane as the robot arm applies a vertical displacement with an object from our object dataset. We measure the distances between the centroids of each dot, which we use to calculate the maximum in-plane strain along the length and width of the membrane. The calculated maximum strain is used to find the vertical displacement required to impart the desired strain for our experiments: low (10%), medium (60%), high (100%).
@@ V-C Baselines
For the tactile-only and membrane mechanics model baselines, we use an unpatterned IR-opaque silicone membrane with the same physical dimensions as the selectively transmissive membrane. We use the Realsense L515 ToF depth camera to collect data for both of these baselines, since using a ToF depth camera with this sensor system can replicate the Soft Bubble sensor [5] and fulfill the input requirement of ToF depth data for the membrane mechanics model. For the proximity-only baseline, we use the selectively transmissive membrane and depth data from the Realsense L515 ToF camera. The computation times for each approach are in Table II.
@@ V-C2 Proximity-Only
The proximity-only approach exclusively uses the proximity point cloud. This baseline uses a reference point cloud of the inflated membrane with no object-induced deformation. The contact patch is estimated as all points closer to the camera than the reference point cloud.
@@ V-C3 Membrane Mechanics Model
Proximity-Only 0.03
Tactile-Only 0.03
Mechanics Model 0.3
TABLE II: Time to compute one frame using an Intel Core i7-8650U CPU.
@@ VI Evaluation
VI-A Experiment Setup
We measure the accuracy of the estimated contact patch when the membrane is in low (approx. 10%), medium (approx. 60%), and high (approx. 100%+) strain states. We use a dataset of three objects: a stuffed octopus, a Rubiks cube, and a cup. These objects were chosen to represent a diverse set of geometries: edges (Rubiks cube), curves (octopus), and non-convex, separate contacts (cup). Each object is held by a Robotiq gripper mounted to a UR10 robot arm and pushed into the sensor. The vertical displacement of the object in each trial induces the desired strain in the membrane: 4 mm for low strain, 24 mm for medium strain, and 120 mm for high strain. For high strain experiments, the object is pressed off-center to allow the uncontacted region to expand outwards and produce high strains in the membrane.
We compare our results to ground truth and three baselines: tactile-only, proximity-only, and the mechanics model-based method. The metric for geometric accuracy of the contact patch point clouds is the average symmetric root mean squared error (RMSE) of the estimated contact patch point cloud and the ground truth point cloud after manual alignment. For the membrane mechanics model-based method, we crop the estimated contact patch to exclude the outer edges of the sensor interface since the model cannot consistently predict the “bulging” behavior of the membrane, leading to spurious estimated contacts.
@@ VI-A Experiment Setup
Low Strain (approx. 10%) Contact Patch RMSE (mm) $\downarrow$
Algorithm octopus cup cube
Fusion 1.79 4.96 1.34
Proximity-Only 2.51 7.41 1.27
Tactile-Only 3.96 6.74 1.58
Mechanics Model 13.73 Fail Fail
Medium Strain (approx. 60%) Contact Patch RMSE (mm) $\downarrow$
Algorithm octopus cup cube
Fusion 2.64 3.1 1.25
Proximity-Only 4.21 8.35 3.43
Tactile-Only 4.32 6.07 1.75
Mechanics Model 9.95 10.38 6.96
High Strain (approx. 100%) Contact Patch RMSE (mm) $\downarrow$
Algorithm octopus cup cube
Fusion 3.93 2.85 1.57
Proximity-Only 2.08 6.63 4.28
Tactile-Only 4.1 4.2 3.79
Mechanics Model 8.31 5.9 8.86
Overall Average for Strain Range Contact Patch RMSE (mm) $\downarrow$
Algorithm low medium high
Fusion 2.7 2.33 2.78
Proximity-Only 3.73 5.33 4.33
Tactile-Only 4.09 4.05 4.03
Mechanics Model 13.73 9.1 7.69
TABLE III: Results from experiments across low, medium, and high strain ranges. Average symmetric RMSE is used to calculate error.
VI-B Results and Analysis
@@
The contact patch estimations are shown in Figure 4 for a qualitative evaluation of contact patch geometry. The average symmetric RMSE analysis per strain state is shown in Table III. Although our proposed fusion algorithm is simple, it generalizes well across our object dataset and strain range, outperforming all of the baselines on average (Figure 7). Fusing both the tactile and proximity modalities is key to overcoming the shortcomings of each modality: the proximity modality tends to overestimate contact while tactile thresholding tends to underestimate contact.
We highlight the fusion algorithm performance for the cup, which requires segmenting two separate contacts with the base and the handle. For the medium and high strain experiments, the fusion algorithm is the only method capable of accurately distinguishing theses separate contacts.
Generally, the proximity modality overestimates contact because it cannot directly sense contact with the object. Tactile thresholding underestimates contact because there is not one threshold that can be applied across all strain ranges. Additionally, tactile thresholding struggles to distinguish two separate contacts.
There are only two specific cases where a baseline has a lower RMSE than the fused contact patch: the proximity-only estimations in the high-strain octopus contact patch and the low-strain cube patch. With both of these cases, the proximity-only baseline overestimation of the contact patch better captured the ground-truth contact patch geometry than the underestimation from the fusion algorithm. In all other cases, the fusion algorithm consistently produced the most accurate contact patch with the lowest RMSE error.
In some low strain experiments, the membrane mechanics model failed to produce an estimate of the contact patch, because the deformation is within the same magnitude as the simulated depth camera noise and thus indistinguishable. We note that the contact patch geometry estimate is dominated by the coarseness of the mesh.
Figure 5: We demonstrate the proposed fusion algorithm for contact patch estimation in two applications with complex membrane mechanics: A. a varied stiffness membrane and, B. membrane wrinkles, or out-of-plane deformations.
@@ VII-A1 Varied Stiffness Membranes
Active and passive variable stiffness materials [27, 28, 29, 30] can be used for real-time optimization of robot end effector shapes and stiffnesses. Here, we demonstrate the robustness of our sensor’s multimodality and contact patch estimation to variations in the sensing membrane’s stiffness patterns (Fig. 5A). We fabricate a varied stiffness membrane by using two base silicones, Dragonskin 10 Fast and Ecoflex 00-35 Fast (Smooth-On), with different Shore hardnesses (Figure 8). Dragonskin 10 is approximately 8x harder than Ecoflex 00-35. The stiff and soft regions are cured together to form one cohesive varied stiffness membrane.
Here, the advantage of our approach is generalization across different stiffness patterns. The proximity-only and tactile-only baseline methods would require tuning for each membrane pattern due to the variance of inflated shapes. The membrane mechanics model would require generating a new mesh that matches both the stiffness pattern and mechanical properties of the different membrane patterns. We run the medium strain cube experiment with the varied stiffness membrane and find the RMSE is 1.09 mm.
@@ VII-A2 Torque and Shear-Induced Wrinkling
A highly deformable contact interface that tolerates a large range of shear and torque before slipping is key to maintaining stiction with a target object. This feature can be critical in mitigating errors in robot manipulators by reducing the requirement for precise perception and control during a task, or simplifying planning by minimizing regrasping or repositioning the object. However, these torques and shear forces can create wrinkles in the deformable membrane surface, which are very difficult to model even with computationally expensive FEM approaches. In this demonstration, we highlight the advantage of our mechanics-independent, fusion contact patch algorithm to estimate the contact patch of a cup while the membrane is wrinkled (Fig. 5B). The robot arm presses the cup into the surface of the sensor and rotates it 45, creating wrinkles behind the cup handle and along the sides of the cup. The estimated contact patch is shown in Figure 8. In comparison with ground truth, the RMSE of the estimated contact patch is 3.49 mm. The geometry of the contact patch qualitatively matches well with ground truth, accurately capturing the side of the cup handle that is in contact with the membrane despite the wrinkles.
@@ VII-B Closed-Loop Control for Whole-Body Manipulation
We demonstrate the real-time speed and integration of the proximity and tactile fusion algorithm with a closed-loop controller to balance a stack of books (Fig. 6). The robot system, Punyo-1 [31], consists of the visuotactile and proximity sensor placed on the robot’s torso, a Kinova Jaco robot arm on each side of the sensor with soft padding, and a Puget computer. An empty cardboard tray is held by Punyo-1 and pressed into the sensorized torso. As each book is placed on the tray held by the robot, the tray’s center of mass changes and the robot adjusts the roll angle of the tray to balance the books. The angle of the tray of books is sensed with the proximity and visuotactile fusion contact patch algorithm.
The sensor runs at 30 FPS and the contact patch is calculated for each frame. To calculate the angle of the tray, we use principal component analysis (PCA) on the contact patch. We found that PCA of the contact patch gives less noisy estimations of the tray position compared to the proximity-only modality, which has added noise from sensing the books. The estimated angle of the tray is given to the whole-body PD controller at a frequency of 29 Hz. We use two arm motion primitives to adjust the tray angle.